OpenStudy (anonymous):

New carpet is installed in a room measuring 18 ft by 14 ft. Find the area of the room in square yards.

OpenStudy (anonymous):

Area= length x width Area= (18 ft)(14ft) then convert your answer in yards (1 square yard= 9 square ft)

OpenStudy (anonymous):

so what would the answer be

OpenStudy (anonymous):

A. 21 yd2 B. 31 yd2 C. 28 yd2 D. 51 yd2 E. 48 yd2

OpenStudy (anonymous):

Did you follow what I did? what will you get if you multiply the dimensions?

OpenStudy (anonymous):

right you'll get 252 ft^2 now we convert it \(252 ft^2 (\frac{ 1 yd^2}{9 ft^2})\)=? what will you get if you divide 252 by 9?

OpenStudy (anonymous):

i got it :) thanks . mind helping me with a few more ?

OpenStudy (anonymous):

A can of paint will cover 200 ft2. How many cans of paint should be purchased in order to paint a cylinder that has a height of 15 ft and a radius of 6 ft?

OpenStudy (anonymous):

do you know the formula for a cylinder?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think its V=πr2h

OpenStudy (anonymous):

that's the volume, you need to find the surface area :) surface area is \(S.A. = 2 \pi r^2 + 2 \pi rh\) your height is 15 ft and radius is 6 ft substitute it in the formula and tell me what you get

OpenStudy (anonymous):

10

OpenStudy (anonymous):

nope, r=6, h=15 plug this values in the formula which I gave you

OpenStudy (anonymous):

i got 791.68

OpenStudy (anonymous):

right :) to find how many cans of paint needed, you have to divide this area by the area that one paint can cover, which is 200 ft^2

OpenStudy (anonymous):

i did that and got 3.9

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

yeah i got the same thing: 3.958 BUT you have to remember that if you'll round this value , it will be 4.0 therefore it will be 4 cans of paint :)

OpenStudy (anonymous):

A walkway 2 m wide surrounds a rectangular plot of grass. The plot is 40 m long and 25 m wide. What is the area of the walkway?

OpenStudy (anonymous):

i got 1000 first but thats not one of my choices . then i divided 1000 and still cant find my answer . i hate math :(

OpenStudy (anonymous):

 first we have to find the area of plot of grass A=length x width Agrass= (40 x 25)=1000 Area of walkaway: , the dimensions of the walkaway is 2 more than the dimensions of the plot so, length= 40+2= 42 and width= 25 + 2 =27 A = (42 x 27)=1134 but take not this is not the area of the walkaway yet, because the plot is inside this walkaway so we need to subtract the area of the walkaway from the are that we found before A=1134-1000=?